OpenStudy (anonymous):

would 7√x*7√x*7√x*7√x be x^4/7 ?

OpenStudy (james1769):

yud

OpenStudy (james1769):

*yes

OpenStudy (anonymous):

thanks

OpenStudy (jdoe0001):

\(\bf 7\sqrt{x}\cdot 7\sqrt{x}\cdot 7\sqrt{x}\cdot 7\sqrt{x}?\)

OpenStudy (anonymous):

yes

OpenStudy (jdoe0001):

\(\bf 7\sqrt{x}\cdot 7\sqrt{x}\cdot 7\sqrt{x}\cdot 7\sqrt{x} \\ \quad \\ (7\cdot 7\cdot 7\cdot 7)\cdot (\sqrt{x}\cdot \sqrt{x}\cdot \sqrt{x}\cdot \sqrt{x}) \implies ?\)

OpenStudy (anonymous):

\[x^{4/7}\]

OpenStudy (jdoe0001):

hmmm not quite.. is just a straight multiplication pretty much so, you'd multiply the coefficients by the coefficients and the radicands by the radicands so \(\bf 7\sqrt{x}\cdot 7\sqrt{x}\cdot 7\sqrt{x}\cdot 7\sqrt{x} \\ \quad \\ (7\cdot 7\cdot 7\cdot 7)\cdot (\sqrt{x}\cdot \sqrt{x}\cdot \sqrt{x}\cdot \sqrt{x})\implies (7^4)(\sqrt{x^4})\implies ?\)

OpenStudy (jdoe0001):

unless, you meant originally \(\Large \sqrt[7]{x}\cdot \sqrt[7]{x}\cdot \sqrt[7]{x}\cdot \sqrt[7]{x}\)

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

thats what I meant

OpenStudy (jdoe0001):

well, hmmm tis, is not the same thing as before :/

OpenStudy (anonymous):

sorry

OpenStudy (jdoe0001):

\(\Large { a^{\frac{{\color{blue} n}}{{\color{red} m}}} \implies \sqrt[{\color{red} m}]{a^{\color{blue} n}} \qquad \qquad \sqrt[{\color{red} m}]{a^{\color{blue} n}}\implies a^{\frac{{\color{blue} n}}{{\color{red} m}}}\qquad thus \\ \quad \\ \quad \\ \sqrt[7]{x}\cdot \sqrt[7]{x}\cdot \sqrt[7]{x}\cdot \sqrt[7]{x}\implies \sqrt[{\color{red}{ 7}}]{x^{\color{blue}{ 4}}}\implies x^{\frac{{\color{blue}{ 4}}}{{\color{red}{ 7}}}} }\)

OpenStudy (anonymous):

thank you so much

OpenStudy (jdoe0001):

yw